# The lattice packing problem in dimension 9 by Voronoi’s algorithm

### v1

@@ 4.2. Implementation details
We implemented Voronoi’s algorithm in highly-optimized C++ code with the usage of several existing C and C++ libraries. We have used both LRS [Avi15] and PPL [BHZ08] for the dual description of the plain dual description cases without symmetry. For the medium-incidence cases we preferred the low-memory usage of LRS, while PPL was typically a bit faster for the low-incidence cases. We used Traces [MP14, MP22] to compute a canonical form (see section 5.2), FLINT [FLINT] for large integer support and the computation of the determinant and Hermite Normal Form of a matrix, and MurmurHash3 as a fast hashing function [App11].
We stored the list of perfect forms in a binary netCDF [Uni22] file. For each perfect form we store the upper diagonal integer coefficients of the scaled canonical form $Q$, and metadata such as $|\operatorname{Min}(Q)|$, $\lambda_{1}(Q)$, $|\operatorname{Aut}(Q)|$, the number of neighbours, its hash, its parents hash, and if it was marked as treated already. This adds up to a total of $147$ bytes per form or roughly 329329GB in total, which drops to roughly an average of $67$ bytes and 150150GB respectively after compression.
@@
In [DSvW24] an enumeration of another kind of geometrical object was done using a parallelized algorithm. We follow a similar approach and explain here our implementation. Our implementation is parallelized via the Message Passing Interface (MPI), where each process is behaving independent and communicates with other processes via messages. This allows to relatively easily scale to hundreds of processes on multiple nodes.
@@ 5.1. Graphs
On the theoretical side, in a groundbreaking work, Babai [Bab16] showed that one can always solve the graph isomorphism problem in quasi-polynomial time $\exp(\log(n)^{c})$ in the number of vertices. Furthermore, this result was extended to a canonical function for graph isomorphism [Bab19]. More importantly for us, on the practical side, there are many highly optimized implementations of such canonical functions, such as the libraries Nauty, Traces [MP14, MP22] and Bliss [JK25], that can work on graphs with up to thousands of vertices. Typically these implementations are limited to a binary weight function with support $\{0,1\}$, but there exist efficient reductions to the binary case using $n\cdot\sqrt{\log_{2}(|\text{supp}(\omega)|)}$ vertices. See the Nauty manual [MP25] for more details.
@@ 5.2. Lattices and PQF’s
In [DSSV07, vW18] the number of perfect forms enumerated was small enough to use a combination of some equivalence invariants, like the determinant, and pair-wise equivalence checks using the backtracking-search approach by Plesken and Souvignier [PS97]. However, this approach is far from sufficient for the billions of perfect forms we have to handle, as many of them have overlapping invariants.
Motivated by this problem we developed a canonical function for quadratic form equivalence, which was first discussed in this context in [vW18] and later published in [DSHVvW20]. The high-level idea is to reduce the canonical form computation of a PQF to that of a graph. To see how this works let’s consider an automorphism $U\in\operatorname{Aut}(Q)$ and the set of minimal vectors $\operatorname{Min}(Q)$ of a PQF $Q$. By definition we have $\operatorname{Min}(Q)=\operatorname{Min}(U^{t}QU)$, but also $\operatorname{Min}(U^{t}QU)=U^{-1}\cdot\operatorname{Min}(Q)$. Automorphisms thus induce a permutation on the set of minimal vectors. Moreover, this permutation is an isometry w.r.t. the inner product induced by $Q$, i.e.,
@@
One can show that the reverse implication is also true if $\operatorname{Span}_{\mathbb{Z}}(\operatorname{Min}(Q))=\mathbb{Z}^{d}$. In that case the problem of PQF equivalence thus becomes one of graph equivalence. Furthermore, any canonical function for graph equivalence can in turn give a canonical form. From our enumeration it follows that the property $\operatorname{Span}_{\mathbb{Z}}(\operatorname{Min}(Q))=\mathbb{Z}^{d}$ is attained for all but one $9$-dimensional perfect lattice and thus using that set is almost always sufficient. We check for this condition and otherwise one has to consider a set of slightly larger vectors. For more details we refer to [DSHVvW20]. A similar approach has later been used in [DSvW24] to build a canonical function for C-type domains.
To further improve the canonical function from [DSHVvW20] we also consider the absolute trick introduced in [vW23, DSvW24]. The idea is that any minimal vector $x\in\operatorname{Min}(Q)$ always comes in a pair $\pm x$. We therefore build an absolute graph with only half the original vertices, typically corresponding to the minimal vectors up to sign $\pm x\in\operatorname{Min}(Q)/\{\pm\}$, and by replacing the weight function by $w(i,j)=|v_{i}^{t}Qv_{j}|$ for $\pm v_{i},\pm v_{j}\in\operatorname{Min}(Q)/\{\pm\}$. The absolute graph is half the size of the original graph and thus we can expect the canonical graph algorithms to be more efficient.
The main problem is that an isomorphism between absolute graphs does not always induce one between the corresponding PQF’s. One can however check this given the automorphisms of the absolute graph that one gets for free from the canonical form computation. If each of them lifts correctly to an automorphism of the PQF we still obtain a canonical form. If not one can still run the computation with the full graph, although this is rarely needed for the perfect forms we consider. For more details see [vW23, Chapter 9.7.2]. When using the Traces library for computing the canonical function on a graph one obtains a speed-up of roughly a factor $3$ from the absolute trick as shown in table 2. In the same table we also see that the computation of the absolute canonical form is only between a factor $1.1$ and $1.6$ times slower than the computation of the automorphism group using the program AUTO by Plesken and Souvignier [PS97], which is comparable to doing a single isomorphism check.
Table 2. Average timing of regular and absolute canonical quadratic form computation over all 99 dimensional perfect forms versus the computation of the automorphism group using the AUTO program [PS97]. The second row displays the weighted average by (orbit) number of neighbours which represents the ratio in which these forms are encountered during Voronoi’s algorithm. The benchmarks were performed on a Xeon Gold SKL-6130 CPU at 2.1 Ghz.
@@
Lastly, we used the results by Watson [Wat71] which claim that the $9$-dimensional perfect forms with $2\cdot 129$ and $2\cdot 136$ minimal vectors respectively are unique, to quickly differentiate those.
@@ 5.3. Polytopes
As combinatorial automorphisms and equivalence are typically hard to compute we will only consider restricted linear equivalence. For this we have a group of linear transformations $G=\operatorname{GL}_{n}(\mathbb{R})$ acting on sets $\{v_{1},\ldots,v_{m}\}\subset\mathbb{R}^{n}$ of full-rank ray generators. Just as for the equivalence of PQF’s we consider a reduction to the case of weighted graphs following [BDSP+14]. In this work it is shown that if one computes the symmetric matrix $S=\sum_{i=1}^{m}v_{i}v_{i}^{t}$, and constructs an $m$-vertex graph with weight function $\omega(i,j)=v_{i}^{t}S^{-1}v_{j}$, that then restricted linear automorphisms and isomorphisms correspond to the respective automorphisms and isomorphisms on these graphs.
@@ 5.4. Permutation groups and face equivalence
The computer algebra system GAP [GAP24] implements many permutation groups algorithm which are very efficient and useful. In particular the partition backtracking search [Leo91, Leo97] is very useful for testing equivalence of faces, and for computing their stabilizer under the symmetry group of a polytope. This was used in [DSSV09, DSSV10, DDS16, BDSS09] for computing the dual description of a variety of polytopes, and in particular for completing Voronoi’s algorithm in dimension $8$ [DSSV07]. For this work, working in GAP was not efficient enough and difficult to integrate with the rest of our C++ implementations. Therefore, the GAP permutation group algorithms that were relevant to the dual description problem were translated to C++ in a new library Permutalib [DS24]. This implementation was already successfully used in [Dut22].
In this work we have to determine the equivalence of up to billions of faces and thus again require a canonical function. Because each orbit is finite one could in principle use the minimal orbit element under some predetermined ordering as a canonical representative. An example of this is to use the standard lexicographic ordering. However, the search for the minimal element under such an ordering can be very costly, depending on the properties of the group and the orbit. One can improve this by (incrementally) determining a suitable ordering depending on the orbit in question, minimizing the backtracking search in the algorithm. Such a canonical function was developed in [JJPW19] and implemented as a GAP package. A generalization of the canonical function code from this GAP package was also translated to Permutalib [DS24].
@@
The group $H$ is used to limit the backtracking search tree in the algorithm, but has to be fixed a priori in a canonical way. In our context, we have two possibilities for the subgroup $H$: the stabilizer $\operatorname{Stab}(G,S)$ of $S$ under $G$ (the choice in [JJPW19]) or the trivial subgroup. The corresponding canonical functions can be different. If the stabilizer is large and one uses the trivial subgroup we could be exposed to memory and runtime problems. If one uses $\operatorname{Stab}(G,S)$ then this problem can be addressed but the computation of the stabilizer is expensive.
@@ 6.1. Recursive Adjacency Decomposition Method
To compute the dual description under symmetries we use the Recursive Adjacency Decomposition Method. The Adjacency Decomposition Method (ADM) has been reinvented many times for example by Jaquet [JC93] for the treatment of $\operatorname{Vor}(Q_{\mathsf{E}_{7}})$ in Voronoi’s algorithm and by Christof and Reinelt [CR96] for cut polytopes, traveling salesman polytopes and the linear ordering polytopes. The algorithm and its name were formally introduced in [CR01].
@@
An important step in the algorithm is the determination of all adjacent facets $F_{1},\ldots,F_{N}$ of a facet $F$. These are in a one-to-one correspondence with the ridges $R_{1},\ldots,R_{N}$ of $P$ for which $R_{i}=F_{i}\cap F$. Given any ridge $R_{i}$ included in $F$ one can uniquely determine $F_{i}$ in a process called flipping which we will detail in section 6.3.2. We can thus focus on obtaining the ridges $R_{i}$ of $P$ which are included in the facet $F$. When viewing the facet $F$ as a subcone (restricted to its linear span), the ridges $R_{1},\ldots,R_{N}$ are precisely the facets of $F$. The determination of the ridges $R_{1},\ldots,R_{N}$ is thus again a dual description problem, but of the subcone $F$ which is in a dimension lower, and given by only those extreme rays of $P$ that are contained in $F$. This typically leads to much easier instances that can be solved by standard dual description algorithms. See the face lattice in fig. 3 which indicates the adjacent facets and the subcone defined by a facet. In several cases these subcones might still be hard to process, therefore [DSSV07] extended the method to a recursive one. On a closer inspection, it is sufficient to enumerate the facet orbits of $F$ up to the symmetries induced by $\operatorname{Stab}(G,F)\subset G$, i.e., any facets $R_{i}$ of $F$ in the same orbit under $\operatorname{Stab}(G,F)$ will lead to an adjacent facet of $P$ in the same orbit under $G$. This is thus again a dual description problem under symmetry which we can solve using ADM. This recursive approach allowed the treatment of the Voronoi domain $\operatorname{Vor}(Q_{\mathsf{E}_{8}})$ of the $\mathsf{E}_{8}$ root lattice in about a month of computation, and thereby led to the successful classification of all perfect forms in dimension $8$.
@@
Algorithm 2 Recursive Adjacency Decomposition Method [CR01, DSSV07]
@@
[…]
@@ 6.2.1. Canonical face representative
Just as in Voronoi’s algorithm one needs to differentiate efficiently between up to billions of facet orbits. The idea of using a canonical function for this was already introduced in [CR01]. In [vW18] it was suggested to use the improved canonical function function from [JJPW19]. We implemented this canonical function, with further improvements as detailed in section 5.4, and used it in our Recursive ADM algorithm.
@@ 6.2.2. Early termination criterion
Theorem 4 (Connectivity theorem [Bal61]).
@@
Therefore, if the set of facets in untreated orbits has size at most n2n-2 then the untreated orbits do not need to be treated since they would not give anything new. This strategy has first been used in [DSV05]. If one inspects the proof of Balinski’s theorem then one sees that only the dimension is used. I.e., given a set of facets and their facet-defining normal vectors $y_{1},\ldots,y_{N}$, if these vectors span a subspace of rank at most n2n-2, then the removal of those facets leaves the graph connected. Therefore if the facets in the untreated orbits have this property we can terminate the enumeration.
Looking even closer into the proof of theorem 4 given by Grünbaum [Grü67], there is a slightly stronger statement in case the set of facets contain a common extreme ray. In the context of the dual description problem this was mentioned in [DDS16, Theorem 1]: if a set of facets all contain the same extreme ray, then removing them leaves the adjacency graph connected.

### v2

@@ 4.2. Implementation details
We implemented Voronoi’s algorithm in highly-optimized C++ code with the usage of several existing C and C++ libraries. We have used both LRS [lrs] and PPL [BagnaraHZ08SCP] for the dual description of the plain dual description cases without symmetry. For the medium-incidence cases we preferred the low-memory usage of LRS, while PPL was typically a bit faster for the low-incidence cases. We used Traces [mckay2014practical, nauty] to compute a canonical form (see section 5.2), FLINT [flint] for large integer support and the computation of the determinant and Hermite Normal Form of a matrix, and MurmurHash3 as a fast hashing function [Appleby_MurmurHash3].
We stored the list of perfect forms in a binary netCDF [NetCDF_Software] file. For each perfect form we store the upper diagonal integer coefficients of the scaled canonical form $Q$, and metadata such as $|\operatorname{Min}(Q)|$, $\lambda_{1}(Q)$, $|\operatorname{Aut}(Q)|$, the number of neighbours, its hash, its parents hash, and if it was marked as treated already. This adds up to a total of $147$ bytes per form or roughly 329329GB in total, which drops to roughly an average of $67$ bytes and 150150GB respectively after compression.
@@
In [IsoEdgeSixDim] an enumeration of another kind of geometrical object was done using a parallelized algorithm. We follow a similar approach and explain here our implementation. Our implementation is parallelized via the Message Passing Interface (MPI), where each process is behaving independent and communicates with other processes via messages. This allows to relatively easily scale to hundreds of processes on multiple nodes.
@@ 5.1. Graphs
On the theoretical side, in a groundbreaking work, Babai [babai2016graph] showed that one can always solve the graph isomorphism problem in quasi-polynomial time $\exp(\log(n)^{c})$ in the number of vertices. Furthermore, this result was extended to a canonical function for graph isomorphism [babai2019canonical]. More importantly for us, on the practical side, there are many highly optimized implementations of such canonical functions, such as the libraries Nauty, Traces [mckay2014practical, nauty] and Bliss [bliss], that can work on graphs with up to thousands of vertices. Typically these implementations are limited to a binary weight function with support $\{0,1\}$, but there exist efficient reductions to the binary case using $n\cdot\sqrt{\log_{2}(|\text{supp}(\omega)|)}$ vertices. See the Nauty manual [mckay2025nauty] for more details.
@@ 5.2. Lattices and PQF’s
In [PerfectDim8, vanWoerdenMasterThesis] the number of perfect forms enumerated was small enough to use a combination of some equivalence invariants, like the determinant, and pair-wise equivalence checks using the backtracking-search approach by Plesken and Souvignier [PleskenSouvignier]. However, this approach is far from sufficient for the billions of perfect forms we have to handle, as many of them have overlapping invariants.
Motivated by this problem we developed a canonical function for quadratic form equivalence, which was first discussed in this context in [vanWoerdenMasterThesis] and later published in [CanonicalFormPositiveForm]. The high-level idea is to reduce the canonical form computation of a PQF to that of a graph. To see how this works let’s consider an automorphism $U\in\operatorname{Aut}(Q)$ and the set of minimal vectors $\operatorname{Min}(Q)$ of a PQF $Q$. By definition we have $\operatorname{Min}(Q)=\operatorname{Min}(U^{t}QU)$, but also $\operatorname{Min}(U^{t}QU)=U^{-1}\cdot\operatorname{Min}(Q)$. Automorphisms thus induce a permutation on the set of minimal vectors. Moreover, this permutation is an isometry w.r.t. the inner product induced by $Q$, i.e.,
@@
One can show that the reverse implication is also true if $\operatorname{Span}_{\mathbb{Z}}(\operatorname{Min}(Q))=\mathbb{Z}^{d}$. In that case the problem of PQF equivalence thus becomes one of graph equivalence. Furthermore, any canonical function for graph equivalence can in turn give a canonical form. From our enumeration it follows that the property $\operatorname{Span}_{\mathbb{Z}}(\operatorname{Min}(Q))=\mathbb{Z}^{d}$ is attained for all but one $9$-dimensional perfect lattice and thus using that set is almost always sufficient. We check for this condition and otherwise one has to consider a set of slightly larger vectors. For more details we refer to [CanonicalFormPositiveForm]. A similar approach has later been used in [IsoEdgeSixDim] to build a canonical function for C-type domains.
To further improve the canonical function from [CanonicalFormPositiveForm] we also consider the absolute trick introduced in [van2023lattice, IsoEdgeSixDim]. The idea is that any minimal vector $x\in\operatorname{Min}(Q)$ always comes in a pair $\pm x$. We therefore build an absolute graph with only half the original vertices, typically corresponding to the minimal vectors up to sign $\pm x\in\operatorname{Min}(Q)/\{\pm\}$, and by replacing the weight function by $w(i,j)=|v_{i}^{t}Qv_{j}|$ for $\pm v_{i},\pm v_{j}\in\operatorname{Min}(Q)/\{\pm\}$. The absolute graph is half the size of the original graph and thus we can expect the canonical graph algorithms to be more efficient.
The main problem is that an isomorphism between absolute graphs does not always induce one between the corresponding PQF’s. One can however check this given the automorphisms of the absolute graph that one gets for free from the canonical form computation. If each of them lifts correctly to an automorphism of the PQF we still obtain a canonical form. If not one can still run the computation with the full graph, although this is rarely needed for the perfect forms we consider. For more details see [van2023lattice, Chapter 9.7.2]. When using the Traces library for computing the canonical function on a graph one obtains a speed-up of roughly a factor $3$ from the absolute trick as shown in table 2. In the same table we also see that the computation of the absolute canonical form is only between a factor $1.1$ and $1.6$ times slower than the computation of the automorphism group using the program AUTO by Plesken and Souvignier [PleskenSouvignier], which is comparable to doing a single isomorphism check.
Table 2. Average timing of regular and absolute canonical quadratic form computation over all 99 dimensional perfect forms versus the computation of the automorphism group using the AUTO program [PleskenSouvignier]. The second row displays the weighted average by (orbit) number of neighbours which represents the ratio in which these forms are encountered during Voronoi’s algorithm. The benchmarks were performed on a Xeon Gold SKL-6130 CPU at 2.1 Ghz.
@@
Lastly, we used the results by Watson [G1971] which claim that the $9$-dimensional perfect forms with $2\cdot 129$ and $2\cdot 136$ minimal vectors respectively are unique, to quickly differentiate those.
@@ 5.3. Polytopes
As combinatorial automorphisms and equivalence are typically hard to compute we will only consider restricted linear equivalence. For this we have a group of linear transformations $G=\operatorname{GL}_{n}(\mathbb{R})$ acting on sets $\{v_{1},\ldots,v_{m}\}\subset\mathbb{R}^{n}$ of full-rank ray generators. Just as for the equivalence of PQF’s we consider a reduction to the case of weighted graphs following [GroupPolytopeLMS]. In this work it is shown that if one computes the symmetric matrix $S=\sum_{i=1}^{m}v_{i}v_{i}^{t}$, and constructs an $m$-vertex graph with weight function $\omega(i,j)=v_{i}^{t}S^{-1}v_{j}$, that then restricted linear automorphisms and isomorphisms correspond to the respective automorphisms and isomorphisms on these graphs.
@@ 5.4. Permutation groups and face equivalence
The computer algebra system GAP [GAP] implements many permutation groups algorithm which are very efficient and useful. In particular the partition backtracking search [Leon1, Leon2] is very useful for testing equivalence of faces, and for computing their stabilizer under the symmetry group of a polytope. This was used in [ComplexityVoronoiDSV, ContactLeech, CUT8_facet, montreal] for computing the dual description of a variety of polytopes, and in particular for completing Voronoi’s algorithm in dimension $8$ [PerfectDim8]. For this work, working in GAP was not efficient enough and difficult to integrate with the rest of our C++ implementations. Therefore, the GAP permutation group algorithms that were relevant to the dual description problem were translated to C++ in a new library Permutalib [permutalib]. This implementation was already successfully used in [BirkhoffDualDesc].
In this work we have to determine the equivalence of up to billions of faces and thus again require a canonical function. Because each orbit is finite one could in principle use the minimal orbit element under some predetermined ordering as a canonical representative. An example of this is to use the standard lexicographic ordering. However, the search for the minimal element under such an ordering can be very costly, depending on the properties of the group and the orbit. One can improve this by (incrementally) determining a suitable ordering depending on the orbit in question, minimizing the backtracking search in the algorithm. Such a canonical function was developed in [jefferson2019minimal] and implemented as a GAP package. A generalization of the canonical function code from this GAP package was also translated to Permutalib [permutalib].
@@
The group $H$ is used to limit the backtracking search tree in the algorithm, but has to be fixed a priori in a canonical way. In our context, we have two possibilities for the subgroup $H$: the stabilizer $\operatorname{Stab}(G,S)$ of $S$ under $G$ (the choice in [jefferson2019minimal]) or the trivial subgroup. The corresponding canonical functions can be different. If the stabilizer is large and one uses the trivial subgroup we could be exposed to memory and runtime problems. If one uses $\operatorname{Stab}(G,S)$ then this problem can be addressed but the computation of the stabilizer is expensive.
@@ 6.1. Recursive Adjacency Decomposition Method
To compute the dual description under symmetries we use the Recursive Adjacency Decomposition Method. The Adjacency Decomposition Method (ADM) has been reinvented many times for example by Jaquet [jaquet1993enumeration] for the treatment of $\operatorname{Vor}(Q_{\mathsf{E}_{7}})$ in Voronoi’s algorithm and by Christof and Reinelt [CR_small_polytopes] for cut polytopes, traveling salesman polytopes and the linear ordering polytopes. The algorithm and its name were formally introduced in [CR_decomposition_parallelization].
@@
An important step in the algorithm is the determination of all adjacent facets $F_{1},\ldots,F_{N}$ of a facet $F$. These are in a one-to-one correspondence with the ridges $R_{1},\ldots,R_{N}$ of $P$ for which $R_{i}=F_{i}\cap F$. Given any ridge $R_{i}$ included in $F$ one can uniquely determine $F_{i}$ in a process called flipping which we will detail in section 6.3.2. We can thus focus on obtaining the ridges $R_{i}$ of $P$ which are included in the facet $F$. When viewing the facet $F$ as a subcone (restricted to its linear span), the ridges $R_{1},\ldots,R_{N}$ are precisely the facets of $F$. The determination of the ridges $R_{1},\ldots,R_{N}$ is thus again a dual description problem, but of the subcone $F$ which is in a dimension lower, and given by only those extreme rays of $P$ that are contained in $F$. This typically leads to much easier instances that can be solved by standard dual description algorithms. See the face lattice in fig. 3 which indicates the adjacent facets and the subcone defined by a facet. In several cases these subcones might still be hard to process, therefore [PerfectDim8] extended the method to a recursive one. On a closer inspection, it is sufficient to enumerate the facet orbits of $F$ up to the symmetries induced by $\operatorname{Stab}(G,F)\subset G$, i.e., any facets $R_{i}$ of $F$ in the same orbit under $\operatorname{Stab}(G,F)$ will lead to an adjacent facet of $P$ in the same orbit under $G$. This is thus again a dual description problem under symmetry which we can solve using ADM. This recursive approach allowed the treatment of the Voronoi domain $\operatorname{Vor}(Q_{\mathsf{E}_{8}})$ of the $\mathsf{E}_{8}$ root lattice in about a month of computation, and thereby led to the successful classification of all perfect forms in dimension $8$.
@@
Algorithm 2 Recursive Adjacency Decomposition Method [CR_decomposition_parallelization, PerfectDim8]
1: A V-representation $V\in\mathbb{Q}^{n\times m}$ of a pointed polyhedral cone $P$, and a group $G\subset\operatorname{Comb}(P)\subset\operatorname{Sym}(m)$.
@@
3: $\operatorname{\mathcal{F}}\leftarrow\{F_{\text{start}}\}$ $\triangleright$ with $F_{\text{start}}$ any facet of $P$.
4: while $\exists$ untreated $F\in L$ do
5:   Compute the facets $R_{1},\ldots,R_{N}$ of $F\subset\operatorname{Span}_{\mathbb{R}}(F)$. $\triangleright$ direct or recurse
6:   for all $R_{i}$ (up to $\operatorname{Stab}(G,F)$) do $\triangleright$ ridges of $P$ included in $F$
7:    Compute unique facet $F_{i}$ of $P$ such that $R_{i}=F\cap F_{i}$. $\triangleright$ flipping
@@
9:      $\operatorname{\mathcal{F}}\leftarrow\operatorname{\mathcal{F}}\cup\{F_{i}\}$.
10:   Mark $F\in\operatorname{\mathcal{F}}$ as treated.
@@ 6.2.1. Canonical face representative
Just as in Voronoi’s algorithm one needs to differentiate efficiently between up to billions of facet orbits. The idea of using a canonical function for this was already introduced in [CR_decomposition_parallelization]. In [vanWoerdenMasterThesis] it was suggested to use the improved canonical function function from [jefferson2019minimal]. We implemented this canonical function, with further improvements as detailed in section 5.4, and used it in our Recursive ADM algorithm.
@@ 6.2.2. Early termination criterion
Theorem 4 (Connectivity theorem [BalinskiTh]).
@@
Therefore, if the set of facets in untreated orbits has size at most n2n-2 then the untreated orbits do not need to be treated since they would not give anything new. This strategy has first been used in [rigid_dim_six]. If one inspects the proof of Balinski’s theorem then one sees that only the dimension is used. I.e., given a set of facets and their facet-defining normal vectors $y_{1},\ldots,y_{N}$, if these vectors span a subspace of rank at most n2n-2, then the removal of those facets leaves the graph connected. Therefore if the facets in the untreated orbits have this property we can terminate the enumeration.
Looking even closer into the proof of theorem 4 given by Grünbaum [grunbaum1967convex], there is a slightly stronger statement in case the set of facets contain a common extreme ray. In the context of the dual description problem this was mentioned in [CUT8_facet, Theorem 1]: if a set of facets all contain the same extreme ray, then removing them leaves the adjacency graph connected.